I have 50.0 grams of SO3. a) how many molecules of SO3 are in this sample? b) how many oxygen atoms are in this sample?
1
Expert's answer
2018-04-02T10:32:05-0400
m(SO3) = 50 g

n(SO3) = m/M = 50 g / 80 g/mol = 0.625 mol

N(SO3) = n*NA = 0.625 mol*6.02*10^23mol-1 = 3.7625*10^23 molecules

N(O) = 3*3.7625*10^23 = 11.2875*10^23 atoms